OpenStudy (anonymous):

The number of kilograms of corn needed to 5000 chickens is 30 less than twice the number of kilograms needed to feed 2800 chickens. How many kilograms of corn are needed to feed 2800 chickens?

OpenStudy (anonymous):

x = 2(c) - 30 x = 2(c) - 30 / 5000 x = 2(c) - 30 / 2800 (5000 / 2800)c 2c - (22 / 28)c = 30 (6 / 28)c = 30 c = 140 140 kilograms of corn are needed to feed 2800 chickens.